OpenStudy (anonymous):

Which expression is equivalent to sec2xcot2x?

OpenStudy (anonymous):

A. sin^2x B. csc^2x C. 1/cos^2x D. 1/sin^2x

OpenStudy (michele_laino):

we have to use these identities: \[\large \begin{gathered} \sec x = \frac{1}{{\cos x}} \hfill \\ \cot x = \frac{{\cos x}}{{\sin x}} \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

ok

OpenStudy (michele_laino):

you should get this: \[\large {\left( {\sec x} \right)^2}{\left( {\cot x} \right)^2} = \frac{1}{{{{\left( {\cos x} \right)}^2}}}\frac{{{{\left( {\cos x} \right)}^2}}}{{{{\left( {\sin x} \right)}^2}}} = ...?\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@matthew1286

OpenStudy (welshfella):

note - the (cos x)^2 cancel out

OpenStudy (anonymous):

ok, well this still dosent make senese

OpenStudy (welshfella):

why ?

OpenStudy (anonymous):

i have no idea

OpenStudy (welshfella):

the question is a bit confusing because sec2x could mean 2 different things - either sec^2x ( which also can be written as (sec x)^2 ) or sec 2x

OpenStudy (anonymous):

^ it has that

OpenStudy (welshfella):

Ok the Michelle's answer is correct you just have to cancel out the (cos x)^2 to give you one of the options

OpenStudy (anonymous):

ok thanks

OpenStudy (welshfella):

yw just remove the 2 (cos x)^2 and see what you have left